OpenStudy (anonymous):

need help on Geometry Part II, exam 4 for ashworth college? does anyone have the answer's for this exam? please let me know asap thank you.

OpenStudy (anonymous):

what is your question

OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1000 earthworms. How many earthworms can he have if his square of compost has a side length that is 8 times longer? A. 8000 B. 64,000 C. 256,000 D. 32,000

OpenStudy (anonymous):

Find the area of a sector with a central angle of 170° and a diameter of 9.1 cm. Round to the nearest tenth. A. 122.9 cm2 B. 30.7 cm2 C. 8.6 cm2 D. 3.4 cm2

OpenStudy (anonymous):

B.\[\frac{170}{360}\pi \left(\frac{9.1}{2}\right)^2=30.713\]

OpenStudy (anonymous):

Hiram raises earthworms. In a square of compost 4 ft by 4 ft, he can have 1000 earthworms. How many earthworms can he have if his square of compost has a side length that is 8 times longer? A. 8000 B. 64,000 C. 256,000 D. 32,000 PROPORTIONS \[\frac{ 1000 }{ 16 } = \frac{ x }{ 1024 }\] find x (the answer is there)

OpenStudy (anonymous):

Thanks.